# Towards Fast Inference: Exploring and Improving Blockwise Parallel Drafts

### v1

@@ Abstract
Despite the remarkable strides made by autoregressive language models, their potential is often hampered by the slow inference speeds inherent in sequential token generation. Blockwise parallel decoding (BPD) was proposed by Stern et al. (2018) as a way to improve inference speed of language models. In this paper, we make two contributions to understanding and improving BPD drafts. We first offer an analysis of the token distributions produced by the BPD prediction heads. Secondly, we use this analysis to inform algorithms to improve BPD inference speed by refining the BPD drafts using small n-gram or neural language models. We empirically show that these refined BPD drafts yield a higher average verified prefix length across tasks.
Language Model
@@ 1 Introduction
The landscape of natural language processing has been profoundly reshaped by recent advances in autoregressive language models (Brown et al., 2020; Wei et al., 2022; Radford et al., 2019; Raffel et al., 2020; Team et al., 2023). These models have shown remarkable proficiency across a range of text generation tasks, including applications like question answering (Rajpurkar et al., 2016) and summarization (Hermann et al., 2015). However, a significant obstacle to their wider application is high inference latency, particularly for extremely deep models with hundreds of billions of parameters (Hoffmann et al., 2022; Rae et al., 2021; Chowdhery et al., 2022). This latency, intrinsic to decoding with autoregressive language models (LMs), imposes considerable computational burdens and limits real-time deployment.
In response to these latency challenges, the field has seen a shift towards decoding methods aimed at reducing the inference latency in large language models (LLM). One promising development is the concept of blockwise parallel decoding (BPD) (Stern et al., 2018; Monea et al., 2023; Cai et al., 2024). Unlike autoregressive decoding, which generates one token at a time, BPD LMs are outfitted with a set of prediction heads, which propose and verify a draft, a block of subsequent tokens, in parallel. While BPD offers one solution to accelerated text generation, it also poses a challenge in ensuring that the proposed drafts are fluent and natural.
Figure 1: Illustration of two tokens that are decoded by autoregressive decoding vs. two tokens drafted by BPD.
BPD inference speed is a function of the time required to produce a BPD draft, and how well that draft matches the decode produced by the base LM (Figure 1). While standard autoregressive LMs generate tokens sequentially, ensuring that each generated token is consistent with all of the preceding tokens (e.g., ‘Messi’ is consistent with ‘Lionel’), BPD adopts a different strategy. At each time step, BPD LMs predict a draft of multiple tokens concurrently (‘Lionel’ and ‘Ronaldo’), each of which is predicted independently of the others. The inherent challenge in BPD drafts lies in maintaining consistency between tokens at each position during parallel generation. A good BPD drafter should prioritize congruous sequences, like ‘Lionel Messi’, over incongruous sequences, like ‘Lionel Ronaldo’, which are unlikely to be decoded by a strong LM. The focus of this work is improving the quality of BPD drafts, without adding additional model parameters that would need to be trained.
@@ 2 Our Contributions
In this paper, we first investigate properties made by the prediction heads of BPD LMs across several tasks; given these observations, we propose rescoring algorithms to produce higher quality BPD drafts.
2.1 Observations on BPD Drafts
Consecutive repetitions  All heads within a block make predictions independently in a BPD LM. Unsurprisingly, we observe that this leads to BPD drafts with significant token repetition across heads. Consecutive repetition is pervasive across tasks, ranging from 20% to 75% of all neighboring draft tokens, depending on the task (subsection 6.1).
Confidence of different heads  We analyze the distribution of probabilities within each softmax head. Our empirical analysis reveals a key property of BPD: the BPD drafter tends to be more confident with initial tokens, and becomes progressively less confident for subsequent tokens. We find that the confidence of BPD heads correlates strongly with the quality of the BPD drafter (subsection 6.2).
Oracle top-$k$ block efficiency  In the BPD algorithm (Algorithm 1), the most likely token at each head is generated as the draft. As highlighted before, this approach is prone to two issues: (1) there might be consecutive repetitions and (2) the model might not be confident about the prediction at some of the heads. We use block efficiency, the average number of draft tokens accepted during decoding, to measure the quality of a given drafter (Leviathan et al., 2023; Sun et al., 2023b). We ask if the block efficiency can be improved by considering the top-$k$ most likely tokens at each head. To measure the potential benefit of considering top-$k$ tokens, we measure the block efficiency of the oracle path through this top-$k$ lattice, oracle top-$k$ block efficiency, and show that there is significant headroom for improvement across tasks (subsection 6.3).
@@ 2.2 New Algorithms
Based on these observations, we propose two algorithms to leverage the top-$k$ predictions at each head and improve BPD latency. Neither of these algorithms require changes to the underlying BPD LM.
Local rescoring via neural model  Given the top-$k$ predictions at each head, we refine the BPD draft by using a small neural, autoregressive LM to greedily rescore these local predictions (subsection 7.1). While the BPD block prediction scores are produced independent of each other, neural rescoring should favor sequences that are fluent, encouraging coherence between the predictions at each head.
Parallel BPD via global n-gram rescoring   If the BPD model has $h$ heads and we consider the top-$k$ tokens from each head, then there are $k^{h}$ candidate drafts of length $h$ that can be formed. We propose to use an n-gram model to efficiently rescore all paths, using dynamic programming, and generate the $p$ most probable rescored paths as a batch of draft candidates. This draft can then be verified in parallel by the BPD model (subsection 7.2).
There are two critical distinctions between the proposed algorithms: the amount of context/expressive power available to each class of rescoring models, and fundamental limitations of decoding with each class. While neural rescoring models are potentially more expressive and can leverage unbounded context, n-gram  LMs can be used to efficiently find the globally most likely rescored drafts from the exponentially-sized set of possible draft candidates. Description of each algorithm is given in section 7.
@@ 2.3 Organization
The remainder of this paper organized as follows. In section 3, we discuss previous literature in reducing LLM latency. In section 4, we define foundational concepts and terminology. section 5 describes our experimental setup, datasets, on methods. section 6 describes our analysis of the BPD drafts. In section 7, we present the proposed BPD rescoring algorithms and empirical results, followed by a final discussion in section 8.
@@ 3.1 Efficient Transformer Inference
Works on improving transformer efficiency encompass both optimization of an existing set of model weights, or a fundamental change to the model architecture. Examples of the former include techniques such as quantization (Xiao et al., 2023; Yao et al., 2022; Dettmers et al., 2022) and model pruning (Sun et al., 2023a; Ma et al., 2023). In parallel, neural architecture search has played a crucial role in identifying network structures that balance performance with efficiency (Kitaev et al., 2020; Zhu et al., 2021). Relatedly, Elbayad et al. (2019) propose early-exiting at intermediate layers for faster inference, while Schuster et al. (2022) explore confidence thresholding for balancing speed and accuracy. These methods offer insights into optimizing decoding under resource constraints.
One important line of work has focused on modifying the decoding method in LMs. The adoption of non-autoregressive (parallel) decoding strategies (Stern et al., 2018; Gu et al., 2017) marks a pivotal shift in this domain, addressing inference latency by simultaneously generating multiple tokens. Subsequent innovations have sought to refine this approach by incorporating additional context (Chi et al., 2020), iterative refinement (Kim et al., 2023), and tree-based attention mechanism (Cai et al., 2024). However, these refinements often require complex training or additional inference data.
@@ 3.2 Efficient and Effective Decoding
There are several recent works that improve the speed of LLM decoding, including pioneering works like BPD and speculative decoding. Speculative decoding leverages a smaller ‘draft’ model to anticipate the outputs of a larger target model, improving average decode latency without loss in generation quality (Leviathan et al., 2023; Chen et al., 2023; Kim et al., 2023). The draft model is typically trained on the same corpus as the LLM, thus autoregressively generates similar drafts as the target model with reduced latency. Speculative decoding is most successful when a long sequence of speculated tokens are accepted by the target LM during verification, avoiding multiple serial calls to the target LM to generate the same sequence.
On the surface, contrastive decoding algorithms share some similarities with our proposed draft rescoring approach, insofar as a weaker model is used to modify the predictions of the target LM (Li et al., 2022; Kim et al., 2024). However, in this work, we refine BPD drafts solely to improve latency. Like speculative decoding, our proposals have no effect on the quality of the target LM’s generated text.
@@ 4 Preliminaries
Autoregressive decoding
Let $\mathcal{M}_{\theta}$ be an autoregressive LM parameterized by $\theta$. The objective is to generate an output sequence $y_{\leq T}=(y_{1},\ldots,y_{T})$ conditioned on an input sequence $\bar{x}$. $z_{t}=\mathcal{M}_{\theta}(y_{t}|\bar{x},y_{<t})$ is a vector of logits, $z_{t}\in\mathbb{R}^{|\mathcal{V}|}$, where $\mathcal{V}$ is the vocabulary over tokens. These logits define a conditional probability distribution at each time step $p_{\theta}(y_{t+1}|\bar{x},y_{\leq t})=\frac{e^{z_{t}}}{\sum e^{z_{t}}}$, which by the chain rule yields $p_{\theta}(y_{\leq T}|\bar{x})=\prod_{t=1}^{T}p_{\theta}(y_{t}|\bar{x},y_{<t})$.
Sequences are generated autoregressively, either through ancestral sampling from some form of the conditional next token distribution (Holtzman et al., 2019), or by a beam search through the space of possible sequences to return a probable sequence. Greedy decoding, a special case of beam search, generates each token as $\hat{y}_{t+1}=\operatorname*{arg\,max}p_{\theta}(y_{t+1}|\bar{x},y_{\leq t})$. In this work, we consider greedy decoding exclusively, as this is the setting that Stern et al. (2018) was designed to accelerate.
Algorithm 1 Blockwise Parallel Decoding (BPD)
0:  : Blockwise parallel decoder $\mathcal{M}_{\theta}^{h}$, initial prompt sequence $\bar{x}$ and target sequence length $T$.
1:  Initialize $t\leftarrow 1$
@@
5:     $y_{t+1}\leftarrow\operatorname*{arg\,max}(z^{1}_{t,1})$
6:     $\hat{y}_{t+2},\dots,\hat{y}_{t+h}\leftarrow$   $\operatorname*{arg\,max}(z^{2}_{t,2}),\cdots,\operatorname*{arg\,max}(z^{h}_{t,h})$
7:     /* Stage 2: Verify */
8:     for $j\leftarrow 2,\dots,h$ in parallel do
9:        $z_{t,j}^{1},\cdots\leftarrow$  $\mathcal{M}_{\theta}^{h}(y_{t+j},\dots|\bar{x},y_{\leq{t+1}},\hat{y}_{t+2},\cdots,\hat{y}_{t+j-1})$
10:     end for
11:     /* Stage 3: Accept */
12:     $n\leftarrow\max\{n:\hat{y}_{t+j}=\operatorname*{arg\,max}z^{1}_{t,j},2\leq j\leq n\}$
13:     $t\leftarrow t+n$
14:  end while
Blockwise parallel decoding
Let $\mathcal{M}^{h}_{\theta}$ be a BPD LM with block size $h$. This model employs $h$ distinct feedforward neural (FFN) layer with a single hidden layer, atop the target LM’s final hidden layer. The output of each FFN is followed by a softmax layer over the vocabulary to predict each of the $h$ subsequent tokens in the block. In our experiments, the parameters of the FFNs are learned jointly with the base LM during training, and the weights of all softmax layers are tied to the input embedding table. Algorithm 1 describes the BPD greedy decoding procedure:
@@
    Verify: At this stage, the target LM greedily generates next-token logits $\{z_{t,2}^{1},\cdots,z_{t,h}^{1}\}$ conditioned on the existing prefix and BPD draft $\{\bar{x},y_{\leq t+1},\hat{y}_{t+2},\cdots,\hat{y}_{t+h}\}$. Verification amounts to checking which BPD draft tokens match the autoregressive greedy decode from the target LM: ($\operatorname*{arg\,max}{z^{1}_{t,i}})==\hat{y}_{t+i}$. Verification of all positions can be performed in parallel under the assumption that the target LM is a decoder-only transformer.
@@
    Accept: Finally, the length of the longest contiguous prefix of draft tokens that match the target LM greedy decode is identified: $n$. The decoded sequence is extended by $n+1$ tokens and we iterate. [^1] The decoded sequence is extended by $n+1$ tokens since during verification we generate the token from the target LM, $\operatorname*{arg\,max}{z_{t,n+1}}^{1}$, at the first position where the draft differs from the target LM greedy decode. Note that in general, not all $h$ tokens are accepted, and many of the draft tokens in each block are discarded. Since the additional time required to generate a block of tokens is fast relative to the time it takes for the forward pass of the target LM, a modest gain in accepted prefix length justifies the cost of draft generation.
@@ 5 Experimental Setup
Table 1: Per-task test performance of each finetuned model and block efficiency. Tasks include Language Modeling (LM), extractive Question Answering (QA), and both Long and Short Summarization (L-Sum & S-Sum). The performance metric for LM is perplexity, for QA is exact match, and for the remaining summarization tasks, the metric is ROUGE-L.
@@
LM LAMBADA (Paperno et al., 2016) 7.88 3.12
QA SQuAD V1 (Rajpurkar et al., 2016) 57.60 2.08
S-SUM CNN/Daily (Hermann et al., 2015) 39.85 1.74
SAMSUM (Gliwa et al., 2019) 37.66 1.27
L-SUM MultiNews (Fabbri et al., 2019) 23.08 1.10
XSUM (Narayan et al., 2018) 52.15 1.13
NewsRoom (Grusky et al., 2018) 39.85 1.08
In this paper, we use $\approx 1.5$ billion (B) parameter decoder-only transformer LMs with up to nine blockwise heads. [^2] This study is based on the original BPD framework, with a modification: we use decoder-only models instead of the T5 encoder-decoder architecture. Other setups are consistent with the approach outlined in Stern et al. (2018). The 1.5B model and all other LMs were pretrained using the C4 English mixture (Raffel et al., 2019) with the causal next token prediction objective tokenized with the GPT3 subword vocabulary (Brown et al., 2020). For the 1.5B BPD models, all heads were trained jointly to predict the following $h$ tokens at each iteration. During pretraining we use batches of 2048 subword sequences, each 512 tokens in length, amounting to $\approx 200$ billion input tokens in total.
We evaluate the potential latency improvement of BPD drafts by block efficiency (Leviathan et al., 2023; Sun et al., 2023b). In this context, block efficiency represents the theoretical speedup compared to standard greedy decoding. It is defined as the average number of tokens decoded per serial call to the BPD model. The formula for block efficiency is
$B\coloneqq\frac{\text{Total number of decoded tokens}}{\text{Total number of serial calls to }\mathcal{M}_{\theta}^{h}}$
In addition to a standard language modeling dataset, LAMBADA (Paperno et al., 2016), we conduct experiments across several classes of downstream tasks. In the realm of text summarization, we evaluate models on the XSUM (Narayan et al., 2018), MultiNews (Fabbri et al., 2019), SAMSum (Gliwa et al., 2019), NewsRoom (Grusky et al., 2018) and CNN/DailyMail (Hermann et al., 2015) datasets. Each of these datasets is characterized by distinct summary lengths and styles. For extractive QA, the SQuAD V1 dataset (Rajpurkar et al., 2016) serves as our testbed. For each task aside from language modeling, we finetune the BPD model for that task. [^3] Details are given in Appendix C.
Table 1 shows that block efficiency varies dramatically across tasks and as a function of the number of BPD heads. We use all nine BPD heads for subsequent experiments as this acts as an upper bound on possible block efficiency.
Table 2: An illustration of outputs from BPD Frameworks on LAMBADA, SQuAD V1, and CNN/Daily datasets. ‘Black’ indicates standard decoded output, ‘Blue’ indicates draft tokens that were accepted during BPD, and ‘Brown’ is the prompt text.

### v2

# Taehyeon Kim1, Ananda Theertha Suresh2 Kishore Papineni2 Michael Riley2 Sanjiv Kumar2 Adrian Benton2 1KAIST AI  2Google Research

@@ Abstract
Despite the remarkable strides made by autoregressive language models, their potential is often hampered by the slow inference speeds inherent in sequential token generation. Blockwise parallel decoding (BPD) was proposed by Stern et al. [38] as a method to improve inference speed of language models by simultaneously predicting multiple future tokens, termed block drafts, which are subsequently verified and conditionally accepted by the autoregressive model. This paper contributes to the understanding and improvement of block drafts in two ways. First, we analyze the token distributions produced by multiple prediction heads. Secondly, we leverage this analysis to develop algorithms to improve BPD inference speed by refining the block drafts using n-gram and neural language models. Experiments demonstrate that refined block drafts yield a +5-21% increase in block efficiency (i.e., the number of accepted tokens from the block draft) across diverse datasets.
@@ 1 Introduction
The landscape of natural language processing has been profoundly reshaped by recent advances in autoregressive language models [3, 43, 30, 33, 42]. These models have shown remarkable proficiency across a range of text generation tasks, including applications like question answering [34] and summarization [15]. However, a significant obstacle to their wider application is high inference latency, particularly for extremely deep models with hundreds of billions of parameters [16, 31, 7]. This latency, intrinsic to decoding with autoregressive language models (LMs), imposes considerable computational burdens and limits real-time deployment.
In response to these latency challenges, the field has seen a shift towards decoding methods aimed at reducing the inference latency in large language models (LLM). One promising development is the concept of blockwise parallel decoding (BPD) [38, 27, 4]. Unlike autoregressive decoding, which generates one token at a time, blockwise parallel LMs are outfitted with a set of prediction heads that propose and verify a draft, a block of subsequent tokens, in parallel. While BPD offers one solution to accelerated text generation, it also poses a challenge in ensuring that the proposed drafts are fluent and natural.
(a) Example of block drafts
(b) Output of our proposed algorithms
Figure 1: (a) Illustration of two tokens that are decoded by autoregressive decoding vs. two tokens drafted by BPD. (b) Outputs from our proposed algorithms, where the top-$k$ token-level predictions are refined using local neural and global n-gram rescoring, which selects the $p$ most probable sequences by dynamic programming, for batched verification.
BPD inference speed depends on both the time it takes to produce a block draft and the draft’s agreement with the base LM’s output (1(a)). Unlike standard autoregressive LMs that generate tokens sequentially—ensuring consistency with all preceding tokens (e.g., ‘Messi’ following ‘Lionel’)—BPD employs a parallel strategy. Here, blockwise parallel LMs simultaneously predict multiple token drafts (e.g., ‘Lionel’ and ‘Ronaldo’), each independently. The primary challenge in BPD is ensuring that these concurrently generated tokens maintain consistency. Effective block drafters should prioritize coherent sequences such as ‘Lionel Messi’ over less likely combinations like ‘Lionel Ronaldo’, which a robust LM would not decode. The focus of this research is improving the quality of block drafts without altering the underlying model parameters.
@@ 2 Our contributions
In this paper, we first investigate properties made by the prediction heads of blockwise parallel LMs across several tasks; given these observations, we propose rescoring algorithms to produce higher quality block drafts.
2.1 Observations on block drafts
Consecutive repetitions  All heads within a block make predictions independently in a blockwise parallel LM. Unsurprisingly, we observe that this leads to block drafts with significant token repetition across heads. Consecutive repetition is pervasive across tasks, ranging from 20% to 75% of all neighboring draft tokens, depending on the task (subsection 6.1).
Confidence of different heads  We analyze the distribution of probabilities within each softmax head. Our empirical analysis reveals a key property of BPD: the block drafter tends to be more confident with initial tokens, and becomes progressively less confident for subsequent tokens. We find that the confidence of block heads correlates strongly with the quality of the block drafter (subsection 6.2).
Oracle top-$k$ block efficiency  In the standard BPD algorithm (Algorithm 1), the most likely token at each head is generated as the draft. As highlighted before, this approach is prone to two issues: (1) there might be consecutive repetitions and (2) the model might not be confident about the prediction at some of the heads. We use block efficiency, the average number of draft tokens accepted during decoding, to measure the quality of a given drafter [24, 41]. We ask if the block efficiency can be improved by considering the top-$k$ most likely tokens at each head. To measure the potential benefit of considering top-$k$ tokens, we measure the block efficiency of the oracle path through this top-$k$ lattice, oracle top-$k$ block efficiency, and show that there is significant headroom for improvement across tasks (subsection 6.3).
@@ 2.2 New algorithms
Based on these observations, we propose two algorithms to leverage the top-$k$ predictions at each head and improve BPD latency (1(b)). Neither of these algorithms require changes to the underlying blockwise parallel LMs.
Local rescoring via neural LMs  Given the top-$k$ predictions at each head, we refine the block draft by using a small neural, autoregressive LM to greedily rescore these local predictions (subsection 7.1). While the block prediction scores are produced independent of each other, neural rescoring should favor sequences that are fluent, encouraging coherence between the predictions at each head.
Global rescoring via n-gram LMs with multi-drafts   If the blockwise parallel LM has $h$ heads and we consider the top-$k$ tokens from each head, then there are $k^{h}$ candidate drafts of length $h$ that can be formed. We propose to use an n-gram model to efficiently rescore all paths, via dynamic programming, and generate the $p$ most probable rescored paths as a batch of draft candidates. These $p$ drafts can then be verified in parallel by the blockwise parallel LM (subsection 7.2).
Figure 2: Relative performance of our methods to standard BPD with a 1.5B parameter blockwise parallel LM on NewsRoom dataset [13]. Details are described in Appendix H.
There are two critical distinctions between the proposed algorithms: the amount of context/expressive power available to each class of rescoring model, and fundamental limitations of decoding with each class. While neural rescoring models are potentially more expressive and can leverage unbounded context, n-gram  LMs can be used to efficiently find the globally most likely rescored drafts from the exponentially-sized set of possible draft candidates. Figure 2 shows that our proposed methods enhance block efficiency, with one approach increasing it by up to +21.30%. This same method also optimizes resource usage, reducing key-value (KV) cache I/O by -2.54% and additionally using FLOPs per token by +4.04%. Description of each algorithm is given in subsection 7.1.
@@ 2.3 Organization
The remainder of this paper organized as follows. In section 3, we discuss previous literature in reducing LLM latency. In section 4, we define foundational concepts and terminology. section 5 describes our experimental setup, datasets, on methods. section 6 describes our analysis of the block drafts. In subsection 7.1, we present the proposed BPD rescoring algorithms and empirical results, followed by a final discussion in section 8.
@@ 3.1 Efficient transformer inference
Works on improving transformer efficiency encompass both optimization of an existing set of model weights, or a fundamental change to the model architecture. Examples of the former include techniques such as quantization [44, 45, 9] and model pruning [40, 26]. In parallel, neural architecture search has played a crucial role in identifying network structures that balance performance with efficiency [22, 46]. Relatedly, Elbayad et al. [10] propose early-exiting at intermediate layers for faster inference, while Schuster et al. [35] explore confidence thresholding for balancing speed and accuracy. These methods offer insights into optimizing decoding under resource constraints.
One important line of work has focused on modifying the decoding method in LMs. The adoption of non-autoregressive (parallel) decoding strategies [38, 14] marks a pivotal shift in this domain, addressing inference latency by simultaneously generating multiple tokens. Subsequent innovations have sought to refine this approach by incorporating additional context [6], iterative refinement [20], and tree-based attention mechanism [4]. However, these refinements often require complex training or additional inference data.
@@ 3.2 Efficient and effective decoding
There are several recent works that improve the speed of LLM decoding, including pioneering works like BPD and speculative decoding. Speculative decoding leverages a smaller ‘draft’ model to anticipate the outputs of a larger target model, improving average decode latency without loss in generation quality [24, 5, 20]. The draft model is typically trained on the same corpus as the LLM, thus autoregressively generates similar drafts as the target model with reduced latency. Speculative decoding is most successful when a long sequence of speculated tokens are accepted by the target LM during verification, avoiding multiple serial calls to the target LM to generate the same sequence.
On the surface, contrastive decoding algorithms share some similarities with our proposed draft rescoring approach, insofar as a weaker model is used to modify the predictions of the target LM [25, 21]. However, in this work, we refine block drafts solely to improve latency. Like speculative decoding, our proposals have no effect on the quality of the target LM’s generated text.
@@ 4 Preliminaries
Autoregressive decoding  Let $\mathcal{M}_{\theta}$ be an autoregressive LM parameterized by $\theta$. The objective is to generate an output sequence $y_{\leq T}=(y_{1},\ldots,y_{T})$ conditioned on an input sequence $\bar{x}$. $z_{t}=\mathcal{M}_{\theta}(y_{t}|\bar{x},y_{<t})$ is a vector of logits, $z_{t}\in\mathbb{R}^{|\mathcal{V}|}$, where $\mathcal{V}$ is the vocabulary over tokens. These logits define a conditional probability distribution at each time step $p_{\theta}(y_{t+1}|\bar{x},y_{\leq t})=\frac{e^{z_{t}}}{\sum e^{z_{t}}}$, which by the chain rule yields $p_{\theta}(y_{\leq T}|\bar{x})=\prod_{t=1}^{T}p_{\theta}(y_{t}|\bar{x},y_{<t})$.
Sequences are generated autoregressively, either through ancestral sampling from some form of the conditional next token distribution [17], or by a beam search through the space of possible sequences to return a probable sequence. Greedy decoding, a special case of beam search, generates each token as $\hat{y}_{t+1}=\operatorname*{arg\,max}p_{\theta}(y_{t+1}|\bar{x},y_{\leq t})$. In this work, we consider greedy decoding exclusively, as this is the setting that Stern et al. [38] was designed to accelerate.
10:  : Blockwise parallel LM $\mathcal{M}_{\theta}^{h}$, initial prompt sequence $\bar{x}$ and target sequence length $T$.
21:  Initialize $t\leftarrow 1$
@@
5:     $y_{t+1},\hat{y}_{t+2},\dots,\hat{y}_{t+h}\leftarrow\operatorname*{arg\,max}(z^{1}_{t,1}),\operatorname*{arg\,max}(z^{2}_{t,2}),\cdots,\operatorname*{arg\,max}(z^{h}_{t,h})$
36:     /* Stage 2: Verify */
7:     for $j\leftarrow 2,\dots,h$ in parallel do
8:        $z_{t,j}^{1},\cdots\leftarrow$$\mathcal{M}_{\theta}^{h}(y_{t+j},\dots|\bar{x},y_{\leq{t+1}},\hat{y}_{t+2},\cdots,\hat{y}_{t+j-1})$
9:     end for
410:     /* Stage 3: Accept */
511:     $n\leftarrow\max\{n:\hat{y}_{t+j}=\operatorname*{arg\,max}z^{1}_{t,j},2\leq j\leq n\}$
612:     $t\leftarrow t+n$
13:  end while
Algorithm 1 Blockwise parallel decoding (BPD)
Blockwise parallel decoding  Let $\mathcal{M}^{h}_{\theta}$ be a blockwise parallel LM with block size $h$. This model employs $h$ distinct feedforward neural (FFN) layer with a single hidden layer, atop the target LM’s final hidden layer. The output of each FFN is followed by a softmax layer over the vocabulary to predict each of the $h$ subsequent tokens in the block. In our experiments, the parameters of the FFNs are learned jointly with the base LM during training, and the weights of all softmax layers are tied to the input embedding table. Algorithm 1 describes the BPD greedy decoding procedure:
@@
    Verify: At this stage, the target LM greedily generates next-token logits $\{z_{t,2}^{1},\cdots,z_{t,h}^{1}\}$ conditioned on the existing prefix and block draft $\{\bar{x},y_{\leq t+1},\hat{y}_{t+2},\cdots,\hat{y}_{t+h}\}$. Verification amounts to checking which block draft tokens match the autoregressive greedy decode from the target LM: ($\operatorname*{arg\,max}{z^{1}_{t,i}})==\hat{y}_{t+i}$. Verification of all positions can be performed in parallel under the assumption that the target LM is a decoder-only transformer.
@@
    Accept: Finally, the length of the longest contiguous prefix of draft tokens that match the target LM greedy decode is identified: $n$. The decoded sequence is extended by $n+1$ tokens and we iterate. [^1] The decoded sequence is extended by $n+1$ tokens since during verification we generate the token from the target LM, $\operatorname*{arg\,max}{z^{1}_{t,n+1}}$, at the first position where the draft differs from the target LM greedy decode. Note that in general, not all $h$ tokens are accepted, and many of the draft tokens in each block are discarded. Since the additional time required to generate a block of tokens is fast relative to the time it takes for the forward pass of the target LM, a modest gain in accepted prefix length justifies the cost of draft generation.
@@ 5 Experimental setup
Table 1: Per-task test performance of each finetuned model and block efficiency over language modeling (LM), extractive question answering (QA), and both long and short summarization (L-Sum & S-Sum).
@@
LM LAMBADA [29] 7.88 3.12
QA SQuAD V1 [34] 57.60 2.08
S-SUM CNN/Daily [15] 39.85 1.74
SAMSUM [12] 37.66 1.27
L-SUM MultiNews [11] 23.08 1.10
XSUM [28] 52.15 1.13
NewsRoom [13] 39.85 1.08
Table 2: Outputs from BPD frameworks. Black indicates standard decoded output, blue indicates accepted draft tokens, and brown is the prompt.
@@
In this paper, we use $\approx 1.5$ billion (B) parameter decoder-only transformer LMs with up to 9 blockwise heads. [^2] This study is based on the original BPD framework, with a modification: we use decoder-only models instead of the T5 encoder-decoder architecture. Other setups are consistent with the approach in Stern et al. [38]. The 1.5B model and all other LMs were pretrained on (English) C4 [32] with the causal next token prediction objective tokenized with the GPT3 subword vocabulary [3]. For the 1.5B blockwise parallel LMs, all heads were trained jointly to predict the following $h$ tokens at each iteration. During pretraining, we use batches of 2048 subword sequences, each 512 tokens in length, amounting to $\approx 200$B input tokens in total on TPUv3/TPUv4 [18] with Jax [2].
We evaluate the potential latency improvement of block drafts by block efficiency [24, 41]. In this context, block efficiency represents the theoretical speedup compared to standard greedy decoding. It is defined as the average number of tokens decoded per serial call to the blockwise parallel LMs. The formula for block efficiency is given by $B\coloneqq\frac{\text{Total number of decoded tokens}}{\text{Total number of serial calls to }\mathcal{M}_{\theta}^{h}}$.
In addition to a standard language modeling dataset, LAMBADA [29], we conduct experiments across several classes of downstream tasks. In the realm of text summarization, we evaluate models on the XSUM [28], MultiNews [11], SAMSum [12], NewsRoom [13] and CNN/DailyMail [15] datasets. Each of these datasets is characterized by distinct summary lengths and styles. For extractive QA, the SQuAD V1 dataset [34] serves as our testbed. For each task aside from language modeling, we finetune the blockwise parallel LM for that task. [^3] Details are given in Appendix C. Table 1444The performance metric for LM is perplexity, for QA is exact match, and for the remaining summarization tasks, the metric is ROUGE-L. shows that block efficiency varies dramatically across tasks and as a function of the number of block draft heads. We use all 9 block draft heads for subsequent experiments as this acts as an upper bound on possible block efficiency.